if the velocity of hydrogen molecule is 5x10^4 cm per sec then its de broglie wavelength

Grade:11

4 Answers

Himanshu
7 years ago
Steps to solve this question :-
  1. The mass of hydrogen molecule = 2 a.m.u. so convert it into grams by dividing 2 by avagardo constant (No).
  2. Now, put the values of mass, planck’s constant and velocity in formula :
Wavelength = (Planck’s constant / Mass × Velocity)
where mass is in kg.
velocity is in ms-1.
and planck’s constant = 6.63 × 10-34 Js-1.
Umakant biswal
7 years ago
@ kriti 
de broglie suggested that when the particle are in motion , they exhibit both wave and particle charecter when they need it . 
so, when the moving particle will behave like wave , then the wavelength possessed by it called as de broglie wavelength . 
acc to the formula of it 
LAMBDA = H / MV 
WHERE lambda is the de broglie wavelength , h is the planks constant ( 6.022 * 10^-34 ) and m is the mass 1 mole of hydrogen atom ., 
acc to the calculation , frst u have to calculate the mass of hydrogen atrom by the following formula 
2/6.022810^-23 
and after u get it that will come in grams . 
and then directly apply the formula , to get the answer 
that is 
lambda = 6.022 * 10 ^-34 / 2/6.022*10^-23 * 5 * 10 ^ 4 
and u will get the answer . 
BUT WHEN CALCULATING IT FOR AN ELECTRON THEN USE THE FOLLOWING FORMULA 
LAMBDA = H / ROOT OVER OF 2MEV 
where E is the charges of an electron and v is its velocity / 
HOPE IT CLEARS YOUR DOUBT 
ALL THE BEST ..
baljinder singh
6 years ago
One hydrogen molecule H2 has mas 2 a.m.u
on first we divide it by avagadro no. we can get in grams ,then divide it by 100o to get mass in kg
now we have mass in kg and also velocity in cm per second after converting velocity in m per second
we use de broglie equation i.e. wavelength=planck constant/ mass * veloctity
by putiing values we get answe in metres
after it should be converted in angstrom
so final answer is 4 angstrom
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

The mass of hydrogen molecule = 2 a.m.u. so convert it into grams by dividing 2 by avagardo constant (No).
Now, put the values of mass, planck’s constant and velocity in formula :Wavelength = (Planck’s constant / Mass × Velocity)
where mass is in kg.
velocity is in ms-1.
and planck’s constant = 6.63× 10-34Js-1

Thanks and Regards